OpenStudy (ashley_f97):

Use basic identities to simplify the expression. (cos^2(θ)/sin^2(θ)) + csc θ sin θ csc2θ sec2θ 1 tan2θ

OpenStudy (anonymous):

cos^2 theta / sin^2 theta + csc theta sin theta = (1 - sin^2 theta) / sin^2 theta + sin theta / sin theta = 1 / sin^2 theta - 1 + 1 = 1 / sin^2 theta = csc^2 theta

OpenStudy (ashley_f97):

Thanks :)